OpenStudy (anonymous):

Multiply. (x^2 + 5)(4x^2 - 3)

OpenStudy (anonymous):

foil man

OpenStudy (anonymous):

FOIL FOR YOUR DEAR LIFE

OpenStudy (anonymous):

4x^4+17x^-15

OpenStudy (anonymous):

17x^2

OpenStudy (anonymous):

use the distributive law (a + b) (c + d) = a(c + d) + b(c + d)

OpenStudy (anonymous):

(x^2 + 5)(4x^2 - 3) = x^2(4x^2 - 3) + 5(4x^2 - 3) = 4x^4 - 3x^2 + 20x^2 - 15 = 4x^4 + 17x^2 - 15